The radius of a bowling ball is 0.11 m and its mass is 7.26 kg. What is its density
(a) in kg m−3?
(b) in g cm−3?
1
Expert's answer
2020-12-24T11:42:26-0500

Answer

a) density is given by

"d=\frac{m}{V}=\frac{7.26}{4/3\pi R^3}"

"=\frac{7.26}{1.33\times3.14\times(0.11)^3}"

"=1,305.75" Kg/m3

b) in gram and centimeter

"d=1305.75\times\frac{1000}{1000000}=1.3g/cm^3"